### v1

@@ 4 Discussion
MBBN demonstrated superior performance across diverse datasets, including UK Biobank, ABCD, and ABIDE, underscoring its robustness across different age groups and developmental stages. While UK Biobank includes primarily middle-aged adults, ABCD and ABIDE focus on children and adolescents during critical brain maturation periods. Unlike most functional brain transformer models ([8, 7, 5]), which are rarely validated on heterogeneous cohorts, MBBN effectively handles lifespan variability, suggesting that it captures fundamental principles of brain organization. This enables accurate predictions of clinical and cognitive outcomes across ages and provides insights into developmental trajectories and age-specific neurological markers.
Our findings revealed distinct, frequency-specific connectivity alterations distinguishing individuals with ADHD and ASD from healthy controls. In ADHD, problematic connectivity was observed in regions essential for visual and audio processing, memory formation, and attention regulation with disruptions varying significantly across frequency ranges.
In the high-frequency range, we observed significant reductions in connectivity between the Parahippocampal area 3 and multiple early visual regions, as well as between the Ventromedial visual area 2 and primary/third visual cortices. These findings imply impaired integration between visual information and memory or contextual processing systems, which may be linked to difficulties in sustaining attention, interpreting visual scenes, and integrating incoming stimuli in ADHD ([63, 64]). In the low-frequency range, decreased connectivity emerged between the Parahippocampal area and anterior temporal, perirhinal, and orbitofrontal regions—networks typically involved in episodic memory, semantic integration, and affective regulation ([65, 66]). Additionally, weakened connections between the primary auditory cortex and multimodal association areas point to disruptions in sensory processing ([67]). These alterations suggest that ADHD may be characterized by deficits in slower, long-range communication necessary for high-level cognitive integration and stable internal state representation. Conversely, the ultralow-frequency range showed increased connectivity in ADHD, particularly between the Ventromedial visual area 3 and spatial attention regions (e.g., Lateral intraparietal area ventral, medial area 7A), as well as between emotion-related regions such as Area TF (the posterior subdivision of the Parahippocampal cortex), Area 25 (Subgenual cingulate cortex), and the anterior cingulate cortex (s32). This pattern may reflect excessive internal processing, emotional dysregulation, and inefficient attentional gating ([68]). Together, these findings indicate that ADHD involves a dynamic reconfiguration of functional brain networks across multiple timescales—with reduced integration in fast and intermediate-range communication, and hyperconnectivity in the slowest neural rhythms—potentially contributing to the core symptoms of inattention, emotional lability, and sensory processing anomalies.
Our frequency-specific analysis of the ABIDE dataset for ASD prediction revealed divergent patterns of functional connectivity alterations across spectral bands, underscoring the multi-timescale nature of network disruptions in autism spectrum disorder. In the high-frequency range, the ASD group exhibited markedly increased connectivity, particularly involving Area PGs (Posterior superior parietal lobule ), Area 8 anterior ventral (within the prefrontal cortex), and Area 8Ad (Anterior dorsal prefrontal cortex) with regions implicated in executive control and higher-order social cognition (e.g., Area p32, Area 9 Middle, Dorsal Area 24d). These patterns suggest hyperconnectivity in frontoparietal and cingulo-opercular circuits, potentially reflecting compensatory or maladaptive neural synchronization related to cognitive control and social attention ([69, 70]). In the low-frequency range, increased connectivity was observed between multiple sensory and associative areas (e.g., Area PH, Primary Sensory Cortex, Auditory and ParaBelt Complexes) and Area posterior 9–46 ventral. This may reflect overintegration or inefficient filtering in low-frequency sensory-cognitive loops, contributing to the sensory hypersensitivity and perceptual integration challenges frequently reported in ASD ([71, 72]). Conversely, the ultralow-frequency range showed decreased connectivity in the ASD group compared to controls, notably in connections between Parahippocampal area 1, Area PH, and auditory complexes with Area 7PC (Precuneus) and Area Frontal Opercular 5. These connections are implicated in high-level integrative functions such as spatial awareness, self-referential processing, and multimodal integration. Reduced ultralow-frequency connectivity may reflect a breakdown in global neural coordination and internal state regulation, core deficits commonly implicated in autism ([73, 74]). Together, these findings reveal a frequency-dependent reconfiguration of functional brain networks in ASD, with fronto-executive hyperconnectivity at higher frequencies and reduced long-range integrative coupling at the slowest rhythms—highlighting potential biomarkers and targets for intervention across multiple neurophysiological timescales.
Connectivity patterns observed across frequency bands revealed a transition from local to more widespread disruptions in both ADHD and ASD. In ADHD, high-frequency alterations were primarily local, focusing on visual and memory-related circuits, while ultralow-frequency changes spanned broader fronto-parietal and cingulo-limbic networks. Similarly, ASD showed local hyperconnectivity in high and low-frequency bands, especially in frontal and sensory areas, with broader hypoconnectivity emerging only in the ultralow-frequency range. These results suggest that while both disorders exhibit frequency-specific connectivity alterations, the spatial scale of disruption expands with slower neural oscillations.
In sum, MBBN’s multifaceted and biologically informed approach to modeling spatiotemporal brain dynamics represents a significant advancement over existing transformer-based frameworks. Its comprehensive integration of neuroscientific principles, self-attention mechanisms, and frequency-specific analyses positions MBBN as a promising foundation for more accurate, interpretable, and clinically meaningful neuroscience research.
@@ Appendix B Interpretation of Network Properties
    Contrasting patterns with FC: While FC exhibits its highest modularity in the raw (unfiltered) condition (0.0630), attention-based connectivity shows essentially no modularity in the raw condition (-0.0000). This striking difference suggests that our attention mechanism captures frequency-specific network organization that is lost when analyzing the full frequency spectrum together.
@@ Appendix D Ablations
MBBN without conv 0.916$\pm$0.004 0.844$\pm$0.006 0.975$\pm$0.001 0.920$\pm$0.001
@@ E.1 Perormance comparison
XGBOOST (CC) 0.768$\pm$0.004 0.696$\pm$0.001 0.859$\pm$003 0.768$\pm$0.005
BNT (CC) 0.895$\pm$0.001 0.814$\pm$0.010 0.950$\pm$0.004 0.864$\pm$0.017
BolT (TD) 0.883$\pm$0.001 0.795$\pm$0.001 0.950$\pm$0.000 0.878$\pm$0.001
vanilla BERT (TD) 0.884$\pm$0.006 0.778$\pm$0.027 0.942$\pm$0.014 0.876$\pm$0.006
MBBN (TD) 0.916$\pm$0.004 0.844$\pm$0.006 0.975$\pm$0.001 0.921$\pm$0.006
@@
BNT (CC) 2.276$\pm$0.015 13.989$\pm$0.682 0.536$\pm$0.008 0.849$\pm$0.003
BolT (TD) 2.217$\pm$0.022 12.729$\pm$0.290 0.541$\pm$0.016 0.843$\pm$0.003
vanilla BERT (TD) 2.194$\pm$0.039 12.751$\pm$0.780 0.5147$\pm$0.014 0.847$\pm$0.003
MBBN (TD) 2.210 $\pm$0.041 13.205$\pm$0.956 0.556$\pm$0.025 0.718$\pm$0.028
@@
vanilla BERT (TD) 1.597$\pm$0.015 4.081$\pm$0.094 0.722$\pm$0.005 0.838$\pm$0.036
MBBN (TD) 1.630 $\pm$0.006 4.225$\pm$0.125 0.717$\pm$0.012 0.825$\pm$0.034
@@
XGBOOST (CC) 0.598$\pm$0.010 0.570$\pm$0.005 0.531$\pm$0.020 0.487$\pm$0.006
BNT (CC) 0.634$\pm$0.008 0.606$\pm$0.014 0.575$\pm$0.013 0.567$\pm$0.019
BolT (TD) 0.603$\pm$0.005 0.592$\pm$0.010 0.650$\pm$0.064 0.598$\pm$0.057
Vanilla BERT (TD) 0.609$\pm$0.020 0.592$\pm$0.015 0.583$\pm$0.004 0.542$\pm$0.021
MBBN [from scratch] (TD) 0.598$\pm$0.009 0.589$\pm$0.024 0.654$\pm$0.134 0.521$\pm$0.059
@@
MBBN [from scratch] (TD) 0.614$\pm$0.020 0.610$\pm$0.011 0.800$\pm$0.067 0.563$\pm$0.051
@@ E.2 FLOPs and number of parameters
XGBOOST (CC) 6100 400
BNT (CC) 3,787,250 1,805,681,216
BolT (TD) 9,597,062 4,226,217,120
Vanilla BERT (TD) 10,050,757 23,497,736,096
MBBN (TD) 11,140,695 70,528,086,432
@@ Appendix F Interpretability
Fourth Visual Area ParaHippocampal Area 3 $<0.001$ HC 0.654
Second Visual Area ParaHippocampal Area 3 $<0.001$ HC 0.645
Polar 10p ProStriate Area $<0.001$ HC 0.643
Primary Visual Cortex VentroMedial Visual Area 2 $<0.001$ HC 0.638
Area PFt ParaHippocampal Area 3 $<0.001$ HC 0.636
Third Visual Area VentroMedial Visual Area 2 $<0.001$ HC 0.636
Second Visual Area VentroMedial Visual Area 2 $<0.001$ HC 0.633
Second Visual Area ProStriate Area $<0.001$ HC 0.633
Area Lateral Occipital 2 VentroMedial Visual Area 2 $<0.001$ HC 0.627
Third Visual Area VentroMedial Visual Area 1 $<0.001$ HC 0.627
Area 11l ProStriate Area $<0.001$ HC 0.626
Third Visual Area ParaHippocampal Area 3 $<0.001$ HC 0.623
Primary Visual Cortex ParaHippocampal Area 3 $<0.001$ HC 0.618
Posterior OFC Complex Frontal Opercular Area 3 $<0.001$ HC 0.616
Area 2 ParaHippocampal Area 3 $<0.001$ HC 0.613
Posterior OFC Complex Anterior Agranular Insula Complex $<0.001$ HC 0.612
Third Visual Area ProStriate Area $<0.001$ HC 0.611
Fourth Visual Area VentroMedial Visual Area 2 $<0.001$ HC 0.61
Area posterior 10p Frontal Opercular Area 4 $<0.001$ HC 0.606
Area 47m Frontal Opercular Area 4 $<0.001$ HC 0.606
Posterior OFC Complex Frontal Opercular Area 4 $<0.001$ HC 0.604
Anterior IntraParietal Area VentroMedial Visual Area 1 $<0.001$ HC 0.601
Area 47l (47 lateral) Frontal Opercular Area 4 $<0.001$ HC 0.597
Area 10r Frontal Opercular Area 4 $<0.001$ HC 0.596
Area 10r Frontal Opercular Area 3 $<0.001$ HC 0.589
Area 11l Dorsal Transitional Visual Area $<0.001$ HC 0.589
Primary Visual Cortex VentroMedial Visual Area 1 $<0.001$ HC 0.589
Area TemporoParietoOccipital Junction 1 ProStriate Area $<0.001$ HC 0.587
Second Visual Area VentroMedial Visual Area 1 $<0.001$ HC 0.587
Orbital Frontal Complex Frontal Opercular Area 4 $<0.001$ HC 0.587
Area 13l Anterior Agranular Insula Complex $<0.001$ HC 0.587
Fourth Visual Area VentroMedial Visual Area 1 $<0.001$ HC 0.584
Posterior InferoTemporal complex VentroMedial Visual Area 2 $<0.001$ HC 0.582
Orbital Frontal Complex Anterior Agranular Insula Complex $<0.001$ HC 0.581
Polar 10p Area 23c $<0.001$ HC 0.578
Polar 10p VentroMedial Visual Area 1 $<0.001$ HC 0.578
Fourth Visual Area Frontal Opercular Area 3 $<0.001$ HC 0.577
Area 31pd VentroMedial Visual Area 1 $<0.001$ HC 0.575
Area Lateral Occipital 1 ProStriate Area $<0.001$ HC 0.575
Insular Granular Complex ProStriate Area $<0.001$ HC 0.574
Primary Auditory Cortex ProStriate Area $<0.001$ HC 0.574
Area 5L ParaHippocampal Area 3 $<0.001$ HC 0.572
Area 47m ProStriate Area $<0.001$ HC 0.571
Anterior IntraParietal Area ParaHippocampal Area 3 $<0.001$ HC 0.57
Area Lateral Occipital 1 ParaHippocampal Area 3 $<0.001$ HC 0.569
Area V4t ParaHippocampal Area 3 $<0.001$ HC 0.568
Area dorsal 23 a+b ParaHippocampal Area 3 $<0.001$ HC 0.567
Area 10v Frontal Opercular Area 4 $<0.001$ HC 0.567
Area V3B Anterior Ventral Insular Area $<0.001$ HC 0.565
Posterior InferoTemporal complex ParaHippocampal Area 3 $<0.001$ HC 0.564
Area 31pd ParaHippocampal Area 3 $<0.001$ HC 0.563
Area V3CD ProStriate Area $<0.001$ HC 0.562
Area V3B Anterior Agranular Insula Complex $<0.001$ HC 0.562
Area V3CD PeriSylvian Language Area $<0.001$ HC 0.562
Middle Temporal Area VentroMedial Visual Area 1 $<0.001$ HC 0.561
Medial Belt Complex ProStriate Area $<0.001$ HC 0.561
Fourth Visual Area ProStriate Area $<0.001$ HC 0.561
Area TemporoParietoOccipital Junction 2 VentroMedial Visual Area 1 $<0.001$ HC 0.56
VentroMedial Visual Area 1 VentroMedial Visual Area 2 $<0.001$ HC 0.559
Polar 10p Area V3B $<0.001$ HC 0.559
Area 52 ProStriate Area $<0.001$ HC 0.559
Third Visual Area Ventral Visual Complex $<0.001$ HC 0.557
Area Lateral Occipital 1 PeriSylvian Language Area $<0.001$ HC 0.556
Area 47l (47 lateral) ProStriate Area $<0.001$ HC 0.556
Area 10v Frontal Opercular Area 3 $<0.001$ HC 0.555
Area 10v Area 23c $<0.001$ HC 0.554
Fourth Visual Area PeriSylvian Language Area $<0.001$ HC 0.553
Area 11l Sixth Visual Area $<0.001$ HC 0.553
Area OP4-PV ParaHippocampal Area 3 $<0.001$ HC 0.553
Third Visual Area PeriSylvian Language Area $<0.001$ HC 0.552
Lateral Area 7A ParaHippocampal Area 3 $<0.001$ HC 0.552
Area 10r Middle Insular Area $<0.001$ HC 0.551
Medial Area 7P ProStriate Area $<0.001$ HC 0.55
Area PGs VentroMedial Visual Area 1 $<0.001$ HC 0.55
Area Lateral IntraParietal dorsal Primary Visual Cortex $<0.001$ HC 0.55
Area 47l (47 lateral) Anterior Ventral Insular Area $<0.001$ HC 0.549
Area TF Area TF $<0.001$ HC 0.549
Medial Area 7P ParaHippocampal Area 3 $<0.001$ HC 0.549
Primary Visual Cortex ProStriate Area $<0.001$ HC 0.549
PreCuneus Visual Area VentroMedial Visual Area 1 $<0.001$ HC 0.549
Area 7PC ParaHippocampal Area 3 $<0.001$ HC 0.549
Area Lateral Occipital 1 Anterior Ventral Insular Area $<0.001$ HC 0.548
Area IntraParietal 0 ParaHippocampal Area 3 $<0.001$ HC 0.548
Area Lateral Occipital 2 ParaHippocampal Area 3 $<0.001$ HC 0.547
Area TE1 posterior Dorsal Transitional Visual Area $<0.001$ HC 0.547
Area PGs ProStriate Area $<0.001$ HC 0.545
Area 31a ParaHippocampal Area 3 $<0.001$ HC 0.545
Area 10v VentroMedial Visual Area 1 $<0.001$ HC 0.545
Area 23d ParaHippocampal Area 3 $<0.001$ HC 0.545
Area FST ParaHippocampal Area 3 $<0.001$ HC 0.545
Area TemporoParietoOccipital Junction 3 VentroMedial Visual Area 1 $<0.001$ HC 0.543
Area 13l Frontal Opercular Area 4 $<0.001$ HC 0.543
Area Lateral Occipital 1 VentroMedial Visual Area 1 $<0.001$ HC 0.543
VentroMedial Visual Area 1 VentroMedial Visual Area 1 $<0.001$ HC 0.542
Area V3CD Anterior Ventral Insular Area $<0.001$ HC 0.542
Area V3CD ParaHippocampal Area 3 $<0.001$ HC 0.542
Area Frontal Opercular 5 ProStriate Area $<0.001$ HC 0.541
Area 31p ventral ParaHippocampal Area 3 $<0.001$ HC 0.541
Area Lateral Occipital 3 Frontal Opercular Area 3 $<0.001$ HC 0.541
Polar 10p Dorsal Transitional Visual Area $<0.001$ HC 0.541

### v2

@@ 4 Discussion
Across diverse cohorts—including UK Biobank, ABCD, and ABIDE—MBBN consistently demonstrated strong predictive performance, capturing neural organizational principles across age groups. Unlike many existing models that are validated only on narrow demographics, MBBN successfully generalized across developmental stages, revealing its potential as a lifespan-inclusive neuroimaging tool.
Our frequency-specific analyses uncovered distinct connectivity disruptions in ADHD and ASD. In ADHD, we observed attenuated high-frequency connectivity in fronto-sensorimotor networks, which may underlie core symptoms such as inattention and executive dysfunction ([51, 52]). At lower frequencies, altered sensory and language processing circuits were evident ([53, 54, 55, 56]), and in the ultralow-frequency range, both hypo- and hyperconnectivity patterns suggested inefficient compensatory mechanisms, particularly involving the parahippocampal and cingulate regions ([57, 58, 59, 60, 61]).
A striking finding was the frequency-dependent shift in Area OP2-3-VS (Opercular parts of Areas 2 and 3 and adjacent Ventral Somatosensory / Sylvian region) connectivity in ADHD. This region consistently emerged as a central node across different frequency bands, suggesting it may serve as a dynamic hub that potentially coordinates neural communication based on oscillatory states. Rather than subserving a fixed function, Area OP2-3-VS appears to assume context-dependent roles, possibly regulating cross-frequency information flow or facilitating inter-network switching—mechanisms that could be critical for coherent cognitive function. In ADHD, disruptions in these putative regulatory processes may contribute to frequency-specific connectivity alterations and the disorder’s heterogeneous symptomatology, underscoring the potential importance of spectral dynamics in characterizing functional hubs.
In ASD, high-frequency hubs involving orbitofrontal and somatosensory circuits were prominent, consistent with known deficits in emotional and social processing ([62, 63, 64, 65]). At low frequencies, the posterior frontopolar cortex emerged as a key integrative hub; disruptions here may impair the coordination of memory, decision-making, and multisensory integration essential for goal-directed behavior ([66, 67, 68, 69]). Interestingly, ultralow-frequency analyses revealed heightened connectivity between the temporo-parietal junction and prefrontal regions, possibly reflecting compensatory cognitive strategies or atypical resource allocation during social cognition tasks ([70, 71, 72, 73]).
Collectively, these results illustrate that MBBN not only outperforms prior models in predictive capacity but also provides interpretable biomarkers aligned with known neurodevelopmental patterns. By aligning transformer-based modeling with core neuroscientific principles—particularly the relevance of neural oscillations—MBBN offers a scalable, biologically plausible framework for understanding individualized brain dynamics.
While MBBN offers notable methodological advancements, some limitations warrant consideration. First, our frequency decomposition employs theory-based spectral bands. Although grounded in established neurophysiological knowledge, this approach may not capture finer-grained oscillatory phenomena that deviate from these canonical bands. Future work could explore more adaptive approaches to move beyond theory-based bands, for example, by employing more granular spectral tiling (e.g., via wavelet packet decomposition [74]) or by developing a fully learnable filter-bank layer within the MBBN architecture. This might allow for end-to-end optimization of frequency bands for specific downstream tasks. Second, the communicability-based masking, while effective in promoting network-aware learning, is computationally demanding. To enhance scalability for ever-larger datasets, future iterations could investigate computationally efficient approximations of communicability, such as those based on polynomial expansions or random walks ([40]). Third, the assessment of test-retest reliability for the derived frequency-specific connectivity patterns would be crucial. Establishing such reliability is not only fundamental to evaluating the robustness of the MBBN model itself but also essential for validating the temporal stability of these patterns. This is particularly important for applications involving the tracking of neurodevelopmental trajectories or disease progression, where consistent and dependable measures are paramount ([75]).
Future extensions of MBBN could incorporate multimodal data (e.g., structural MRI, diffusion MRI) and expand into task-based paradigms to examine how frequency-specific connectivity evolves under cognitive demands. Additionally, exploring cross-frequency coupling may offer deeper insights into hierarchical brain dynamics beyond individual frequency bands. MBBN establishes a new standard for interpretable, frequency-resolved brain modeling. By integrating neurophysiological principles with transformer architectures, it offers both predictive power and mechanistic insights into brain function, laying the foundation for precision neuroscience applications.
@@ Appendix B Interpretation of Network Properties
    Contrasting patterns with FC: While FC exhibits its highest modularity in the raw (unfiltered) condition (0.0630), attention-based connectivity shows essentially no modularity in the raw condition ($<0.0001$). This striking difference suggests that our attention mechanism captures frequency-specific network organization that is lost when analyzing the full frequency spectrum together.
@@ Appendix D Ablations
MBBN without conv 0.916$\pm$0.004 0.844$\pm$0.006 0.980$\pm$0.001 0.920$\pm$0.001
@@ E.1 Performance comparison
XGBOOST (CC) 0.785$\pm$0.005 0.709$\pm$0.003 0.859$\pm$003 0.768$\pm$0.005
BNT (CC) 0.872$\pm$0.005 0.795$\pm$0.007 0.950$\pm$0.004 0.864$\pm$0.017
BolT (TD) 0.886$\pm$0.002 0.795$\pm$0.002 0.950$\pm$0.000 0.878$\pm$0.001
vanilla BERT (TD) 0.864$\pm$0.004 0.777$\pm$0.006 0.942$\pm$0.014 0.876$\pm$0.006
MBBN (TD) 0.913$\pm$0.002 0.823$\pm$0.009 0.980$\pm$0.001 0.921$\pm$0.006
@@
BNT (CC) 2.276$\pm$0.015 13.989$\pm$0.682 0.546$\pm$0.015 0.847$\pm$0.006
BolT (TD) 2.217$\pm$0.022 12.729$\pm$0.290 0.588$\pm$0.006 0.847$\pm$0.001
vanilla BERT (TD) 2.194$\pm$0.039 12.751$\pm$0.780 0.546$\pm$0.015 0.818$\pm$0.051
MBBN (TD) 2.210 $\pm$0.041 13.205$\pm$0.956 0.590$\pm$0.013 0.750$\pm$0.008
@@
vanilla BERT (TD) 1.597$\pm$0.015 4.081$\pm$0.094 0.699$\pm$0.016 0.788$\pm$0.036
MBBN (TD) 1.630 $\pm$0.006 4.225$\pm$0.125 0.694$\pm$0.018 0.774$\pm$0.041
@@
XGBOOST (CC) 0.598$\pm$0.010 0.570$\pm$0.005 0.573$\pm$0.053 0.546$\pm$0.046
BNT (CC) 0.634$\pm$0.008 0.606$\pm$0.014 0.691$\pm$0.168 0.617$\pm$0.062
BolT (TD) 0.603$\pm$0.005 0.592$\pm$0.010 0.632$\pm$0.005 0.566$\pm$0.003
Vanilla BERT (TD) 0.609$\pm$0.020 0.592$\pm$0.015 0.644$\pm$0.089 0.625$\pm$0.051
MBBN [from scratch] (TD) 0.633$\pm$0.009 0.593$\pm$0.015 0.780$\pm$0.012 0.657$\pm$0.038
@@
MBBN [from scratch] (TD) 0.614$\pm$0.020 0.610$\pm$0.011 0.800$\pm$0.067 0.683$\pm$0.062
@@ E.2 FLOPs and number of parameters
XGBOOST (CC) 6004 400
BNT (CC) 11,678,810 13,084,233,536
BolT (TD) 20,697,661 50,032,700,640
Vanilla BERT (TD) 22,280,137 69,848,980,256
MBBN (TD) 23,370,075 209,581,818,912
@@ Appendix F Interpretability
Area OP4-PV R Area 10v R HC $<0.001$ 0.564
Primary Motor Cortex R Area 10v R HC $<0.001$ 0.561
Primary Motor Cortex L Area 10v R HC $<0.001$ 0.56
Medial Belt Complex R Area 10v R HC $<0.001$ 0.559
Area OP1-SII R Area 10v R HC $<0.001$ 0.559
Area PF Opercular R Area 10v R HC $<0.001$ 0.558
Area 43 L Area 10v R HC $<0.001$ 0.557
Area a24 R Ventral Area 6 L HC $<0.001$ 0.554
Superior Temporal Visual Area R Area 10v R HC $<0.001$ 0.554
Area Posterior 24 prime R Area 10v R HC $<0.001$ 0.554
Area 1 R Area 10v R HC $<0.001$ 0.553
PreSubiculum R Area 10v R HC $<0.001$ 0.551
Anterior IntraParietal Area R Area 10v R HC $<0.001$ 0.55
Area PFt R Area 10v R HC $<0.001$ 0.549
Area posterior 24 L Ventral Area 6 L HC $<0.001$ 0.548
Ventral Area 6 L Area 10v R HC $<0.001$ 0.547
Area 2 R Area 10v R HC $<0.001$ 0.546
Anterior Ventral Insular Area L Area 10v R HC $<0.001$ 0.545
Area PFcm R Area 10v R HC $<0.001$ 0.544
Frontal Opercular Area 3 L Area 10v R HC $<0.001$ 0.543
Ventral IntraParietal Complex L Area 10v R HC $<0.001$ 0.541
Area PFm Complex R Area 10v R HC $<0.001$ 0.536
VentroMedial Visual Area 1 R Area 10v R HC $<0.001$ 0.536
Area 7PC R Area 10v R HC $<0.001$ 0.536
Area V3B R Area 10v R HC $<0.001$ 0.535
Area posterior 24 L Area 10v R HC $<0.001$ 0.535
RetroInsular Cortex R Area 10v R HC $<0.001$ 0.535
Area TE1 anterior L Ventral Area 6 L HC $<0.001$ 0.534
Ventral Area 24d R Area 10v R HC $<0.001$ 0.534
Area 25 R Ventral Area 6 L HC $<0.001$ 0.534
Area p32 prime R Area 10v R HC $<0.001$ 0.534
Insular Granular Complex R Area 10v R HC $<0.001$ 0.534
Area 43 R Area 10v R HC $<0.001$ 0.531
Area 1 L Area 10v R HC $<0.001$ 0.53
Area TA2 L Area 10v R HC $<0.001$ 0.53
Primary Sensory Cortex R Area 10v R HC $<0.001$ 0.53
Area PGi R Area 10v R HC $<0.001$ 0.53
Para-Insular Area L Area 10v R HC $<0.001$ 0.529
Frontal Eye Fields L Area 10v R HC $<0.001$ 0.529
Area 46 R Area 10v R HC $<0.001$ 0.528
Area OP2-3-VS R Area 10v R HC $<0.001$ 0.527
Area 55b R Area 10v R HC $<0.001$ 0.527
Area dorsal 32 R Ventral Area 6 L HC $<0.001$ 0.527
VentroMedial Visual Area 2 R Area 10v R HC $<0.001$ 0.526
Dorsal Transitional Visual Area R Area 10v R HC $<0.001$ 0.526
Area dorsal 32 L Area 11l R HC $<0.001$ 0.525
Frontal Opercular Area 2 R Area 10v R HC $<0.001$ 0.525
Area OP2-3-VS L Area 10v R HC $<0.001$ 0.524
Area 10r R Ventral Area 6 L HC $<0.001$ 0.524
Area 10v L Ventral Area 6 L HC $<0.001$ 0.524
Frontal Opercular Area 4 L Ventral Area 6 L HC $<0.001$ 0.524
Rostral Area 6 L Area 10v R HC $<0.001$ 0.523
Frontal Opercular Area 1 R Area 10v R HC $<0.001$ 0.523
Area PF Complex R Area 10v R HC $<0.001$ 0.522
Area STSv anterior L Ventral Area 6 L HC $<0.001$ 0.521
Area s32 L Ventral Area 6 L HC $<0.001$ 0.52
Area TemporoParietoOccipital Junction 1 L Area 10v R HC $<0.001$ 0.52
Frontal Opercular Area 4 L Area 10v R HC $<0.001$ 0.52
Area 13l R Area 10v R HC $<0.001$ 0.52
Area 3a R Area 10v R HC $<0.001$ 0.519
Area PGi R Primary Motor Cortex R HC $<0.001$ 0.518
Dorsal Area 24d R Area 10v R HC $<0.001$ 0.518
Area dorsal 32 L Ventral Area 6 L HC $<0.001$ 0.518
Area Posterior 24 prime L Area 10v R HC $<0.001$ 0.518
Medial Belt Complex R Primary Motor Cortex R HC $<0.001$ 0.518
Area posterior 24 R Ventral Area 6 L HC $<0.001$ 0.517
Area PF Opercular L Area 10v R HC $<0.001$ 0.517
Area IntraParietal 2 R Area 10v R HC $<0.001$ 0.517
Area 47m R Area 10v R HC $<0.001$ 0.517
Area 25 L Ventral Area 6 L HC $<0.001$ 0.517
RetroSplenial Complex R Area 10v R HC $<0.001$ 0.516
Parieto-Occipital Sulcus Area 1 L Area 10v R HC $<0.001$ 0.516
Sixth Visual Area R Area 10v R HC $<0.001$ 0.516
Area Frontal Opercular 5 L Area 11l R HC $<0.001$ 0.516
Orbital Frontal Complex L Ventral Area 6 L HC $<0.001$ 0.515
ParaHippocampal Area 3 R Area 10v R HC $<0.001$ 0.515
Area p32 R Ventral Area 6 L HC $<0.001$ 0.514
Area 10d L Ventral Area 6 L HC $<0.001$ 0.514
Area a24 L Ventral Area 6 L HC $<0.001$ 0.514
Area 43 L Ventral Area 6 L HC $<0.001$ 0.514
Frontal Opercular Area 3 R Area 10v R HC $<0.001$ 0.513
Supplementary and Cingulate Eye Field R Area 10v R HC $<0.001$ 0.513
Dorsal area 6 L Area 10v R HC $<0.001$ 0.513
Area PFt L Area 10v R HC $<0.001$ 0.513
Area V6A R Area 10v R HC $<0.001$ 0.513
Insular Granular Complex L Area 10v R HC $<0.001$ 0.513
Area PFcm L Area 10v R HC $<0.001$ 0.513
Anterior 24 prime R Area 10v R HC $<0.001$ 0.513
Anterior Ventral Insular Area L Ventral Area 6 L HC $<0.001$ 0.513
ParaHippocampal Area 1 R Area 10v R HC $<0.001$ 0.512
PeriSylvian Language Area L Area 10v R HC $<0.001$ 0.511
RetroSplenial Complex R Area Posterior Insular 1 L HC $<0.001$ 0.511
Area 9-46d R Area 10v R HC $<0.001$ 0.511
Area posterior 24 L Area 11l R HC $<0.001$ 0.511
VentroMedial Visual Area 3 R Area 10v R HC $<0.001$ 0.511
Area 47s L Area 11l R HC $<0.001$ 0.511
Parieto-Occipital Sulcus Area 2 R Area 10v R HC $<0.001$ 0.51
Area TG dorsal L Area 10v R HC $<0.001$ 0.51
Area 2 L Area 10v R HC $<0.001$ 0.51
Supplementary and Cingulate Eye Field L Area 10v R HC $<0.001$ 0.51